Bill can clean his office in 4 hours. Bob can do the cleaning in 3. How many hours will it take them if they work together?
An inlet pipe can fill an artificial pond in 60 minutes, while an outlet pipe an empty it in 80. Through an error, both pipes are left on. How long will it take for the pond to fill?

Answer by jim_thompson5910(35256)

I'll do the first one to get you started

Let t = time it takes them to do the job together

1/4 + 1/3 = 1/t

3/12 + 4/12 = 1/t

7/12 = 1/t

7*t = 12*1

7t = 12

t = 12/7

t = 1 & 5/7

So it will take exactly 1 & 5/7 hours (roughly 1.714 hours or roughly 1 hour, 42.857 minutes) to do the job together